OpenStudy (anonymous):

f(x) = 2x + 3 g(x) = 2x^2+ 2x Find f(g(x) )

OpenStudy (anonymous):

@jim_thompson5910 same thing just reverse right?

jimthompson5910 (jim_thompson5910):

yep, you start with f(x) now then you replace every x with g(x)

OpenStudy (anonymous):

oh

jimthompson5910 (jim_thompson5910):

f(x) = 2x + 3 f(g(x)) = 2*g(x) + 3 .. replace every x with g(x) f(g(x)) = 2*(2x^2+2x) + 3 ... replace g(x) on the right with 2x^2+2x simplify

OpenStudy (anonymous):

ok so i got 4x^2+4x+3

jimthompson5910 (jim_thompson5910):

same here

OpenStudy (anonymous):

:)